Find the Greatest common factor of the set of monomials
44s^4t^2,84s^2t
Please go step by step on how you got your answer.

Answer by Edwin McCravy(20059)

44s4t2, 84s2t 

Break each down to the product of primes and 
single letter factors:

44s4t2 = 2·2·11·s·s·s·s·t·t

84s2t = 2·2·3·7·s·s·t

Pick all but only the factors that they have in common 
which are two "2"'s, two "s"'s and 1 "t".

GCF = 2·2·s·s·t = 4s2t

Edwin